OpenStudy (marissalovescats):

@jim_thompson5910 Determine whether the Mean Value Theorem applies to f(x)=3x-x^2 on the interval [2,3]. If the Mean Value Theorem applies, find all value(s) of c in the interval such that f'(c) - f(3)-f(2)/3-2. If it does not apply, state why.

jimthompson5910 (jim_thompson5910):

ok, the first condition we need to check is to see if f(x) is continuous on [2,3]

jimthompson5910 (jim_thompson5910):

because if it is not, then the MVT won't work

OpenStudy (marissalovescats):

It is

jimthompson5910 (jim_thompson5910):

yes because it's continuous everywhere (it's a polynomial)

OpenStudy (marissalovescats):

Yup

jimthompson5910 (jim_thompson5910):

also, we can easily determine that f(x) is differentiable everywhere so f(x) is differentiable on (2,3)

jimthompson5910 (jim_thompson5910):

so the MVT applies in this case

jimthompson5910 (jim_thompson5910):

now you need to compute (f(3) - f(2))/(3 - 2)

jimthompson5910 (jim_thompson5910):

what do you get?

OpenStudy (marissalovescats):

I gotttt -2

jimthompson5910 (jim_thompson5910):

me too

OpenStudy (marissalovescats):

I set f'(x) equal to that and solve for x.. right?

jimthompson5910 (jim_thompson5910):

now you need to calculate f ' (x) then set it equal to -2 and solve for x to get the value of c such that f ' (c) = -2

jimthompson5910 (jim_thompson5910):

yes correct

OpenStudy (marissalovescats):

x=5/2

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

so c = 5/2

OpenStudy (marissalovescats):

Now what

jimthompson5910 (jim_thompson5910):

you found the value of c such that \[\Large f ' (c) = \frac{f(3)-f(2)}{3-2}\] so you're done with the problem

jimthompson5910 (jim_thompson5910):

c = 5/2

OpenStudy (marissalovescats):

Oh yaya lol

jimthompson5910 (jim_thompson5910):

and again, the MVT applies because f(x) is continuous on [2,3] AND differentiable on (2,3)

jimthompson5910 (jim_thompson5910):

so that's why it works

OpenStudy (marissalovescats):

I didnt get full credit on this one: Find the minimum and maximum values of f(x)=x^2-2x+1 on the interval [0,3] I got x=1 and yeah I could only find a min at x=1?

jimthompson5910 (jim_thompson5910):

x = 1 is where the min happens, but you need to plug it back into f(x) to get the actual min value

jimthompson5910 (jim_thompson5910):

also, you need to test the endpoints

OpenStudy (marissalovescats):

I did that. He wants a aximum too.

OpenStudy (marissalovescats):

The min is that (1,0)

jimthompson5910 (jim_thompson5910):

with closed intervals, the abs min/max could be at the endpoints

jimthompson5910 (jim_thompson5910):

yes the min is at (1,0) the min value is 0

OpenStudy (marissalovescats):

Min value is 1

jimthompson5910 (jim_thompson5910):

no the min value is 0 because that is the result of f(1)

jimthompson5910 (jim_thompson5910):

now you need to calculate f(0) and f(3)

OpenStudy (marissalovescats):

I dont understand lol

jimthompson5910 (jim_thompson5910):

the endpoints of the interval are at x = 0 and x = 3 this is because the interval is [0,3]

jimthompson5910 (jim_thompson5910):

f(x)=x^2-2x+1 f(0)=0^2-2(0)+1 f(0) = 1

jimthompson5910 (jim_thompson5910):

f(x)=x^2-2x+1 f(3)=(3)^2-2(3)+1 f(3) = 4

jimthompson5910 (jim_thompson5910):

notice how x = 3 produces the largest output f(3) = 4 is the largest you can get

jimthompson5910 (jim_thompson5910):

so the absolute max is at (3,4) and that max value is 4

jimthompson5910 (jim_thompson5910):

where are you stuck?

OpenStudy (marissalovescats):

I got that now

OpenStudy (marissalovescats):

isnt 0 an abs max too then since its a porabola?

jimthompson5910 (jim_thompson5910):

the abs min is 0 because this the lowest point it happens at (1,0)

OpenStudy (marissalovescats):

Oh got it

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):